• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 16:04:21 on June29,2005

    Entry number 147497

    keyword=Shift summary

    10:30 Noticed that averaged energy spread (based on SLI) jumped from around 3.5x10-5 to about 3.7x10-5. Call MCC and ask them to take a look.
    11:10 MCC call to let us know that beam is off due to RF separator problem.
    11:35 MCC called to give us beam back.
    11:55 Call MCC to let them know that SLI signal is still bad (around 5x10-5). They said they are still watching it.
    14:40 Right septum magnet trip.
    15:15 Right septum is back
    15:25 During Run 4995, noticed that there was a shift in beam energy. Stop run and call MCC to fix the energy feedback.




    Run# Start Stop #Events Comments
    4984 07:44 08:33 1M
    4985 08:34 09:19 1M
    4986 09:20 10:08 1M
    4987 10:09 10:57 1M
    4988 10:58 11:05 No beam during run
    4989 11:07 11:08 No beam, RF separator prob
    4990 11:37 12:23 1M High E spread in beginning
    4991 12:23 13:10 1M
    4992 13:11 14:01 1M
    4993 14:01 14:34 0.6M Right septum trip
    4994 DAQ propblem
    4995 15:16 15:22 Change in beam energy, junk
    4996 15:26 On going